the perimeter of a rectangular carpet is 70 feet. the width is three-fourths the length. find the width.
Answer by checkley77(12844)

W=3/4L
P=2L=2W
70=2L+2(3/4L)
70=2L+6L/4
70=2L+1.5L
70=3.5L
L=70/3.5
L=20 FT. ANS. FOR THE LENGTH.
W=3/4*20=15 ANS. FOR THE WIDTH.